OpenStudy (anonymous):

Find the equation of the axis of symmetry. f(x)=-3(x-1)^2-6 PARABOLA

OpenStudy (amistre64):

the equation eh....

OpenStudy (anonymous):

ya...that what i said haha

OpenStudy (amistre64):

y = M(x-k)+h ^ x = k

OpenStudy (anonymous):

u sure?

OpenStudy (amistre64):

yep; unless they changed :)

OpenStudy (anonymous):

haha ok thanks just making sure !! :)

OpenStudy (amistre64):

i gave a general reply; do you know what the answer to the specific problem is tho?

OpenStudy (amistre64):

f(x)=-3(x-1)^2-6 ^ x=1 would be the equation

OpenStudy (anonymous):

no